# Online Slip Detection and Friction Coefficient Estimation for Autonomous Racing

### v2

@@ Abstract
Keywords: Slip detection, friction coefficient estimation, autonomous racing
@@ I INTRODUCTION
The effectiveness of vehicles depends on accurate knowledge of key vehicle and environment parameters, among which the tire-road friction coefficient (TRFC) is one of the most critical [1, 2]. The TRFC defines the upper limit of forces that can be transmitted between tires and road, thereby constraining braking, steering, and overall vehicle stability [3]. Accurate and real-time estimation of this parameter is essential not only for human drivers, but increasingly for automated decision-making, and trajectory planning in autonomous vehicles [4]. In the domain of autonomous racing, operating at or near these friction limits is essential to achieve competitive lap times. Online slip detection is therefore a crucial capability: it provides feedback for exploiting available traction while maintaining stability, and it supports higher-level functions such as adaptive control, trajectory planning, and race strategy. Moreover, since tire conditions strongly influence traction availability, slip detection also plays a key role in monitoring and managing the effective use of tires during high-performance driving [5].
Unlike velocity or acceleration, however, the TRFC cannot be measured directly by standard onboard sensors. Specialized sensors capable of direct measurement exist; however, their high cost and integration challenges prevent their widespread use. This limitation has motivated extensive research into indirect estimation methods [6].
@@
Within effect-based approaches, a large body of research has adopted model-based techniques such as Kalman filtering. Extended Kalman Filters (EKF) [9], Unscented Kalman Filters (UKF) [11], and Cubature Kalman Filters (CKF) [12] have all been applied to TRFC estimation, fusing standard vehicle sensors with simplified models. These approaches can significantly improve estimation accuracy, but they remain model-dependent and are sensitive to parameter uncertainty and nonlinearities.
@@
The contributions of this paper are:
  • A lightweight slip detection technique using only IMU and LiDAR readings and control actions.
  • A real-time estimation method for TRFC without explicit tire/vehicle modeling or training data.
  • Experimental validation demonstrating reasonable online performance of our approach.
@@ II-A Vehicle Level
The plant $P$ is the physical system under consideration, including its geometry and physical parameters. The plant has a center of mass (CoM), where the total mass $m$ is assumed to be concentrated. A right-handed body-fixed 3D coordinate system $B=\{X,Y,Z\}$ is attached at point $CoM$, as shown in Fig. 1:
  • Longitudinal axis $X$ points forward along the chassis
  • Lateral axis $Y$ points leftward at 90° to axis $X$
  • Normal axis $Z$ points upward
In our case, the plant is a 1/10-scale autonomous racing car (also referred to as the vehicle), and it is characterized by a set of configuration, inertial, and geometric properties. The configuration is defined by four tires in contact with the road. Each tire has an effective tire radius $r_{e}$, defined as the distance from the tire center to the contact point as shown in Fig. 2. The vehicle is rear-wheel drive (RWD), while the front axle is actuated (steerable). The inertial properties of the vehicle include its total mass $m$ and the yaw moment of inertia $I_{z}$ about the normal axis $Z$.
The geometric properties of our vehicle, illustrated in Fig. 1, consist of the distances from the CoM to the front and rear axles, denoted by $l_{f}$ and $l_{r}$, respectively.
@@
The time-variant status of plant $P$ is characterized by $n$ state variables $\mathcal{V}$. The state space $\mathcal{S}\subseteq\mathbb{R}^{n}$ is the set of all states (value vectors) $s$ of variables $\mathcal{V}$.
@@
In our case, the observation space $\mathcal{Y}\subseteq\mathbb{R}^{5}$ includes the outputs of the following sensors:
  • Inertial Measurement Unit (IMU) measures CoM’s longitudinal acceleration $\hat{a}_{x}$ and lateral acceleration $\hat{a}_{y}$.
  • LiDAR measures the CoM’s longitudinal velocity $\hat{v}_{x}$, lateral velocity $\hat{v}_{y}$, and the yaw rate $\hat{\omega}_{\psi}$.
Thus, the observation vector $y\in\mathcal{Y}$ is
@@
Our action space $\mathcal{U}\subseteq\mathbb{R}^{2}$ contains a control vector $u\in\mathcal{U}$ equal to $u=\begin{bmatrix}v&\delta\end{bmatrix}^{\top}$, coming from a human or software:
  • Desired longitudinal velocity $v$,
  • Desired steering angle $\delta$.
@@
$\dot{x}$ $=v\cos(\psi+\beta)$
$\dot{y}$ $=v\sin(\psi+\beta)$ (3)
$\dot{\psi}$ $=\frac{v}{l_{r}}\sin(\beta),$
@@
A force $F\in\mathbb{R}^{3}$ is a vector in coordinate system $B$, representing an interaction between bodies, and consists of these components:
  • Longitudinal force $F_{x}$ acts along the forward $X$-axis of the body, positive when pushing forward.
  • Lateral force $F_{y}$ acts along the $Y$-axis, positive when pushing to the left.
  • Normal force $F_{z}$ acts along the $Z$-axis, positive when pushing upward.
@@
Under the bicycle model, subscripts f and r represent the front and rear axles, respectively, the set of forces $F$ consists of tire forces, rolling resistance force $F_{rr}$, and aerodynamical force $F_{drag}$: $F=\{F_{x_{f}},F_{y_{f}},F_{x_{r}},F_{y_{r}},F_{drag}\}$. The vehicle dynamics in the $X_{t}Y_{t}$ plane in $B$ are given by:
  1. 1.
    Longitudinal motion:
@@
In this paper, however, we do not employ any dynamical model.
@@ II-B Tire Level
Given plant $P$ with coordinate system $B=\{X,Y,Z\}$, we attach a local body-fixed 3D coordinate system $B_{w}=\{X_{w},Y_{w},Z_{w}\}$ to the center of each tire. The $X_{w}$ axis points the rotating direction of the tire, the $Y_{w}$ axis points leftward at 90° to the $X_{w}$ axis, and the $Z_{w}$ axis is parallel to $Z$ (see Fig. 2).
@@
  • Slip angle: The slip angle $\alpha$ is between the tire longitudinal axis $X_{w}$ and the tire center velocity vector $V_{w}$ as shown in Fig. 2.
@@
Definition 11 (Friction Coefficient)
@@ II-C Problem Statement
  • No parametric models are used
@@ III RELATED WORK
Cause-based approaches measure friction-related parameters of the road/tire and correlate them with friction in several ways. Tread sensing monitors deformations at the tire–road interface to infer friction [22, 23]. Optical methods use cameras or laser-based sensors to detect surface properties such as texture, water, or ice, sometimes employing infrared or polarization measurements, and neural networks trained on image features [24, 25, 26]. Acoustic methods instead classify the road type from tire noise or microphone signals, often combined with machine learning techniques such as support vector machines or neural networks [27, 28]. While these approaches bypass the need for vehicle or tire models, they require heavyweight hardware instrumentation and are sensitive to environmental disturbances, which limits their deployment in practice. In contrast, our approach relies only on the standard onboard vehicle sensors.
@@
Effect-based approaches can be split into two main categories: model-based and model-free approaches. The former aim to estimate TRFC using simplified mathematical or dynamical representations of the tire–vehicle system, typically without requiring additional external sensors. These approaches are generally more accurate and repeatable than cause-based approaches, and can themselves be broadly divided into three categories [21]. First, vehicle dynamics methods employ simplified dynamical models such as the single-wheel, quarter-car, bicycle, or four-wheel vehicle models, combined with estimation algorithms including recursive least squares, Kalman filters, or sliding mode observers, to infer unmeasured variables such as tire forces, slip ratios, and ultimately the friction coefficient [29, 30]. Second, tire-based methods exploit mathematical tire models, such as the Magic Formula, Dugoff, Brush, or LuGre models, which relate tire forces and moments to slip ratio and slip angle. They estimate the maximum tire–road friction by fitting or observing deviations between measured and modeled behavior [31, 32, 33]. Finally, slip-slope based approaches rely on the relationship between the initial linear region of the $\mu$–slip curve and its saturation level, using the slope at low slip as a predictor of the maximum available friction [17, 34]. Collectively, these methods represent the dominant body of work on TRFC estimation in the literature.
@@
Yet, the effectiveness of model-based approaches is fundamentally tied to accurate knowledge of model parameters and tire–road forces, which are often difficult or impractical to obtain. Tire models, for example, require prior calibration of parameters such as cornering stiffness [9] or Magic Formula coefficients [35, 31], while vehicle dynamical models depend on quantities like mass, moment of inertia, and external forces [35]. Estimating these values typically involves additional filtering or indirect measurements [11], but even then, parameters such as the moment of inertia can vary significantly with changes in mass distribution, leading to degraded performance. This strong dependency on parameter identification limits the adaptability of model-based methods, motivating the need for model-free alternatives that bypass explicit parameterization altogether.
@@
Unfortunately, most model-free approaches depend on extensive datasets or specialized sensing modalities, which limit their adaptability and generalization. In contrast, our work proposes a lightweight formulation that detects slip and estimates friction directly from motion signals, without requiring significant model fitting or large training data.
@@ IV-A1 Linear Detection
Slip along the $X$ axis is identified by comparing the expected velocity $v$, predicted from the commanded throttle by the vehicle’s onboard odometry system, to the measured longitudinal velocity $\hat{v}_{x}$, obtained from sensor fusion of LiDAR and IMU measurements using an EKF for localization. The EKF incorporates a parameter-free, omnidirectional kinematic model to update the state estimates — without relying on explicit dynamical or tire-force modeling. This provides a consistent velocity estimate that can be directly compared to the commanded input.
@@ IV-A2 Angular Detection
The expected angular velocity $\omega_{\psi}$ is derived from the kinematic bicycle model $f$ as follows. From Eq. 3, the yaw rate is given by $\dot{\psi}=\frac{v}{l_{r}}\sin(\beta),$ where $\beta$ is the geometric slip angle. Substituting the expression for $\beta$ from Eq. 4, $\beta=\arctan\!\left(\frac{l_{r}}{l_{f}+l_{r}}\tan(\delta)\right),$ into the yaw rate equation, we obtain
$\dot{\psi}=\frac{v}{l_{r}}\sin\!\left(\arctan\!\left(\frac{l_{r}}{l_{f}+l_{r}}\tan(\delta)\right)\right),$ (10)
which simplifies to
$\dot{\psi}=\frac{v}{l_{f}+l_{r}}\tan(\delta)$ (11)
@@ IV-B Friction Estimation
At the tire level, $\mu$ is defined in terms of the longitudinal and lateral forces acting on a single tire. However, to avoid building a tire model, we combine the forces at the vehicle level. In this view, all tire–road interactions are aggregated into two external forces (longitudinal and lateral) acting at the vehicle’s center of mass: one longitudinal ($F_{x}$) and one lateral ($F_{y}$). The resulting friction coefficient should therefore be interpreted as the average value across all four tires. As such, we assume a negligible discrepancy between the tires.
@@
$F_{x}=m\hat{a}_{x},\quad F_{y}=m\hat{a}_{y}$ (15)
and on flat ground, the normal force is approximated as
$F_{z}=mg$ (16)
@@
Figure 3: Maps of the closed-loop track used for testing.
@@ V Experimental Evaluation
The goals of the experimental evaluation are threefold:
  1. 1.
    Can our approach detect whether a slip occurred?
  2. 2.
    How soon can our approach detect the slip?
  3. 3.
    How accurately can our approach estimate the tire-road friction coefficient over different surfaces?
V-A Experimental Setup
To evaluate the effectiveness of our approach in the autonomous vehicle setting, we conduct experiments on a closed-loop track composed primarily of ceramic tile, with one section embedded with cardboard to provide two surfaces with varying friction. The layout of our track is shown in Figure 3. To ensure reproducibility, each lap began at the marked starting point, from where the vehicle was manually driven counter-clockwise to execute one maneuver to generate slip on each of the two surfaces. We followed this approach 20 times, logging each lap into a ROS2 bag (with corresponding video footage) for analysis and testing. In total, our dataset comprises 20 bags, providing 40 slip events – 20 on each surface.
Our RoboRacer test vehicle [38] is lightweight, both in terms of hardware and computational resources. The onboard NVIDIA Jetson Xavier NX board has limited computing capacity, and we rely on LiDAR and IMU data rather than camera data.
This lightweight platform struggles to run more computationally intensive baselines and lacks instrumentation to build complex dynamical and tire models. Because of this, we focused our evaluation on methods that can run on our platform.
@@ V-B Data Processing
V-B1 Data Labeling
We collected 20 ROS2 bags with the onboard data and the corresponding third-person videos of the vehicle movement. Next, we manually labeled the videos to identify the time of each slip. This allowed us to properly evaluate both the accuracy and delay of our slip detection module. First, the videos were synced to the bags based on the instant the steering angle control stick was released, as it is clearly visible in both the recorded videos and the bag data. Then, the slips were labeled based on when the car visibly began to slide laterally. This time has a downside of potentially being later than the actual slip time, but this labeling procedure allows for consistency.
V-B2 Threshold Cross-validation
@@
To avoid overfitting to specific runs in our evaluation, we perform K-fold cross-validation for the detection thresholds $\delta_{lin}$ and $\delta_{ang}$. We used $K=5$ folds, each consisting of 4 randomly selected bags. In each fold, 16 training bags are used to set the threshold (where the thresholds are two standard deviations above the means of our heuristics $|v-\hat{v}_{x}|$ and $|\omega_{\psi}-\hat{\omega}_{\psi}|$). Then, the remaining 4 test bags were used for evaluation. This was repeated for all 5 folds, ensuring that every bag was used as a testing bag once. This testing data was then used for evaluating slip detection accuracy, delay, and friction estimation accuracy.
V-C Ground-Truth Friction Measurement

### v3

@@ I Introduction
Effective road driving depends on accurate knowledge of key vehicle and environment parameters, among which the tire-road friction coefficient (TRFC) is one of the most critical [20, 21]. The TRFC defines the upper limit of forces that can be transmitted between tires and road, thereby constraining braking, steering, and overall vehicle stability [24]. Accurate and real-time estimation of this parameter is essential not only for human drivers, but increasingly for automated decision-making, including trajectory planning in autonomous vehicles [7]. In autonomous racing, operating at or near these friction limits is essential to achieve competitive lap times. Therefore, online slip detection is a crucial capability: it provides feedback for exploiting available traction while maintaining stability, and it supports higher-level functions such as adaptive control, trajectory planning, and race strategy. Moreover, since tire conditions strongly influence traction availability, slip detection also plays a key role in monitoring and managing the effective use of tires during high-performance driving [16].
Unlike velocity or acceleration, however, the TRFC cannot be measured directly by standard onboard sensors. Specialized sensors capable of direct measurement exist; however, their high cost and integration challenges prevent their widespread use. This limitation has motivated extensive research into indirect TRFC estimation methods [36].
@@
Within effect-based approaches, a large body of research has adopted model-based techniques such as Kalman filtering. Extended Kalman Filters (EKF) [3], Unscented Kalman Filters (UKF) and Model Predictive Control (MPC) [34], and Cubature Kalman Filters (CKF) [38] have all been applied to TRFC estimation, fusing standard vehicle sensors with simplified models. These approaches can significantly improve estimation accuracy, but they remain model-dependent and are sensitive to parameter uncertainty and nonlinearities.
@@
This paper makes three contributions: (i) A lightweight slip detection technique using only IMU and LiDAR readings and control actions; (ii) A real-time estimation method for TRFC without explicit tire/vehicle modeling or training data; (iii) An experimental validation on 1/10-scale racing cars, showing reasonable online performance of our approach.
@@ II-A Vehicle Level
The plant $P$ is the physical system under consideration, including its geometry and physical parameters. The plant has a center of mass (CoM), where the total mass $m$ is assumed to be concentrated. A right-handed body-fixed 3D coordinate system $B=\{X,Y,Z\}$ is attached at point $CoM$, as shown in Figure 1 (Longitudinal axis $X$ points forward along the chassis, lateral axis $Y$ points leftward at 90° to axis $X$, and the normal axis $Z$ points upward).
In our case, the plant is a 1/10-scale autonomous racing car (also referred to as the vehicle), and it is characterized by a set of configuration, inertial, and geometric properties. The configuration is defined by four tires in contact with the road. Each tire has an effective tire radius $r_{e}$, defined as the distance from the tire center to the contact point as shown in Figure 2. The vehicle is rear-wheel drive, while the front axle is actuated (steerable).
The inertial properties of the vehicle include its total mass $m$ and the yaw moment of inertia $I_{z}$ about the normal axis $Z$. The geometric properties of our vehicle, illustrated in Figure 1, consist of the distances from the CoM to the front and rear axles, denoted by $l_{f}$ and $l_{r}$, respectively.
@@
The time-variant status of plant $P$ is characterized by $n$ state variables $\mathcal{V}$. The state space $\mathcal{S}\subseteq\mathbb{R}^{n}$ is the set of all feasible states (value vectors) $s$ of variables $\mathcal{V}$.
@@
In our case, the observation space $\mathcal{Y}\subseteq\mathbb{R}^{5}$ includes the outputs of two sensors: (i) an Inertial Measurement Unit (IMU) measures CoM’s longitudinal acceleration $\hat{a}_{x}$ and lateral acceleration $\hat{a}_{y}$; (ii) a LiDAR measures the CoM’s longitudinal velocity $\hat{v}_{x}$, lateral velocity $\hat{v}_{y}$, and the yaw rate $\hat{\omega}_{\psi}$. The IMU measurements were calibrated to compensate for bias, scale factor errors, and axis misalignment, yielding corrected acceleration estimates used in the state observation. The LiDAR-based velocity estimates were obtained using the standard sensor configuration, without additional calibration adjustments.
Formally, the observation vector $y\in\mathcal{Y}$ is
@@
Our action space $\mathcal{U}\subseteq\mathbb{R}^{2}$ contains a control vector $u\in\mathcal{U}$ equal to $u=\begin{bmatrix}v&\delta\end{bmatrix}^{\top}$, coming from a human or software: desired longitudinal velocity $v$ and desired steering angle $\delta$.
@@
$\dot{x}=v\cos(\psi+\beta),\quad\dot{y}=v\sin(\psi+\beta),\quad\dot{\psi}=\frac{v}{l_{r}}\sin(\beta)$ (3)
@@
A force $F\in\mathbb{R}^{3}$ is a vector in coordinate system $B$, representing an interaction between bodies, and consists of three components $(F_{x},F_{y},F_{z})$, where the longitudinal force $F_{x}$ acts along the forward $X$-axis of the body (positive when pushing forward), the lateral force $F_{y}$ acts along the $Y$-axis (positive when pushing to the left), and the normal force $F_{z}$ acts along the $Z$-axis (positive when pushing upward).
@@
Under the bicycle model, subscripts f and r represent the front and rear axles, respectively, the set of forces $F=\{F_{x_{f}},F_{y_{f}},F_{x_{r}},F_{y_{r}},F_{drag}\}$ consists of tire forces, rolling resistance force $F_{rr}$, and aerodynamical force $F_{drag}$. The vehicle dynamics in the $X_{t}Y_{t}$ plane in $B$ are given by:
@@
This paper, however, does not employ any dynamical model.
@@ II-B Tire Level
Given plant $P$ with coordinate system $B=\{X,Y,Z\}$, we attach a local body-fixed 3D coordinate system $B_{w}=\{X_{w},Y_{w},Z_{w}\}$ to the center of each tire. The $X_{w}$ axis points in the tire’s rotating direction, the $Y_{w}$ axis points leftward at 90° to the $X_{w}$ axis, and the $Z_{w}$ axis is parallel to $Z$ (see Figure 2).
@@
  • Slip angle: The slip angle $\alpha$ is between the tire longitudinal axis $X_{w}$ and the tire center velocity vector $V_{w}$ as shown in Figure 2.
@@
Definition 11 (Tire Road Friction Coefficient)
@@ II-C Problem Statement
  • No parametric models are learned.
@@ III Related Works
Cause-based approaches measure friction-related parameters of the road/tire and correlate them with friction in several ways. Tread sensing monitors deformations at the tire–road interface to infer friction [11] [6]. Optical methods use cameras or laser-based sensors to detect surface properties such as texture, water, or ice, sometimes employing infrared or polarization measurements, and neural networks trained on image features [2] [35]. Acoustic methods instead classify the road type from tire noise or microphone signals, often combined with machine learning techniques such as support vector machines or neural networks [1] [18]. While these approaches bypass the need for vehicle or tire models, they require heavyweight hardware instrumentation and are sensitive to environmental disturbances, which limits their deployment in practice. In contrast, our approach relies only on the standard onboard vehicle sensors.
@@
Effect-based approaches can be split into two main categories: model-based and model-free approaches. The former aim to estimate TRFC using simplified mathematical or dynamical representations of the tire–vehicle system, typically without requiring additional external sensors. These approaches are generally more accurate and repeatable than cause-based approaches, and can themselves be broadly divided into three categories [17]. First, vehicle dynamics methods employ simplified dynamical models such as the single-wheel, quarter-car, bicycle, or four-wheel vehicle models, combined with estimation algorithms including recursive least squares, Kalman filters, or sliding mode observers, to infer unmeasured variables such as tire forces, slip ratios, and ultimately the friction coefficient [40, 39, 19, 28, 15]. Second, tire-based methods exploit mathematical tire models, such as the Magic Formula, Dugoff, Brush, or LuGre models, which relate tire forces and moments to slip ratio and slip angle. They estimate the maximum tire–road friction by fitting or observing deviations between measured and modeled behavior [26, 33, 12]. Finally, slip-slope based approaches rely on the relationship between the initial linear region of the $\mu$–slip curve and its saturation level, using the slope at low slip as a predictor of the maximum available friction [29, 37]. Collectively, these methods represent the dominant body of work on TRFC estimation in the literature.
@@
Yet, the effectiveness of model-based approaches is fundamentally tied to accurate knowledge of model parameters and tire-road forces, which are often difficult or impractical to obtain. Tire models, for example, require prior calibration of parameters such as cornering stiffness [3] or Magic Formula coefficients [8, 26], while vehicle dynamical models depend on quantities like mass, moment of inertia, and external forces [8]. Estimating these values typically involves additional filtering or indirect measurements [34], but even then, parameters such as the moment of inertia can vary significantly with changes in mass distribution, leading to degraded performance. This strong dependency on parameter identification limits the adaptability of model-based methods, motivating the need for model-free alternatives that bypass explicit parameterization altogether.
@@
Unfortunately, most model-free approaches depend on extensive datasets or specialized sensing modalities, which limit their adaptability and generalization [32]. In contrast, our work proposes a lightweight formulation that detects slip and estimates friction directly from motion signals, without requiring significant model fitting or large training data.
@@ IV-A1 Linear Detection
Slip along the $X$ axis is identified by comparing the expected velocity $v$, predicted from the commanded throttle by the vehicle’s onboard odometry system, to the measured longitudinal velocity $\hat{v}_{x}$, obtained from an EKF that fuses IMU measurements with pose estimates from Monte Carlo localization (MCL) [10]. MCL localizes the vehicle within the map using LiDAR, while the EKF incorporates these pose estimates together with IMU data and a parameter-free, omnidirectional kinematic model to provide the full state, including vehicle velocities. This approach provides a velocity estimate that can be directly compared to the commanded input without relying on explicit dynamical or tire-force modeling.
@@ IV-A2 Angular Detection
The expected angular velocity $\omega_{\psi}$ is derived from the kinematic bicycle model $f$ as follows. From Eq. 3, the yaw rate is given by
$\dot{\psi}=\frac{v}{l_{r}}\sin(\beta)$.
Substituting the expression for $\beta$ from Eq. 4
into the yaw rate equation, we obtain
$\dot{\psi}=\frac{v}{l_{r}}\sin\!\left(\arctan\!\left(\frac{l_{r}}{l_{f}+l_{r}}\tan(\delta)\right)\right)=\frac{v}{l_{f}+l_{r}}\tan(\delta).$ (10)
@@ IV-B Friction Estimation
At the tire level, $\mu$ is defined in terms of the longitudinal and lateral forces acting on a single tire. However, to avoid building a tire model, we combine the forces at the vehicle level. Then, all tire-road interactions are aggregated into two external forces (longitudinal and lateral) acting at the vehicle’s center of mass: one longitudinal ($F_{x}$) and one lateral ($F_{y}$). The resulting friction coefficient should therefore be interpreted as the average value across all four tires. As such, we assume a negligible discrepancy between the tires.
@@
$F_{x}=m\hat{a}_{x},\quad F_{y}=m\hat{a}_{y},\quad F_{z}=mg$ (14)
@@
Figure 3: Map of the first, second, and third closed-loop tracks used for testing.
@@ V Experimental Evaluation
We evaluate our approach along three dimensions: (1) slip detection accuracy, (2) slip detection delay, and (3) tire–road friction coefficient (TRFC) stimation accuracy.
Figure 4: Friction circle visualization for cardboard, acrylic, and tile surfaces. The blue dots denote measured longitudinal and lateral accelerations (in g); the star indicates the maximum traction point corresponding to the estimated friction coefficient $\mu$.
V-A Experimental Setup Overview
Experiments were conducted using the RoboRacer platform [25], a lightweight 1:10-scale autonomous racing vehicle equipped with an IMU and LiDAR and powered by an onboard Jetson Xavier NX. The platform does not include camera-based perception or high-fidelity vehicle instrumentation, and therefore represents a resource-constrained embedded setting. An anonymized implementation of our ROS2 drift detector and the data is available online: https://github.com/Trustworthy-Engineered-Autonomy-Lab/Online-Slip-Detection-Friction-Estimation
Testing was performed on three closed-loop tracks composed primarily of ceramic tile with embedded sections of cardboard and acrylic to provide three distinct friction regimes. Across all tracks, 56 recorded runs produced 148 labeled slip events spanning all surface types.
The three tracks differed in their combinations of tile, cardboard, and acrylic sections, providing two-three distinct friction regimes per track. The specifics are shown in Figure 3.
@@ V-B Data Processing
Data Labeling
For each track, ROS2 bags were recorded together with synchronized third-person video. Slip events were manually labeled by aligning the video and bags using the steering-release timestamp and marking the first visible lateral deviation of the vehicle. Although this visual criterion may slightly lag the true physical onset of slip, it provides a consistent and repeatable labeling procedure.
Threshold Cross-validation
To avoid overfitting to specific runs in our evaluation, we perform K-fold cross-validation for the detection thresholds $\delta_{lin}$ and $\delta_{ang}$ using the recorded dataset across all tracks, which consists of 56 valid bags. We used $K=5$ folds. Cross-validation was performed independently on each track to ensure that the detection thresholds were not biased toward specific runs or surface configurations.
@@
In each fold, the training bags were used to set the thresholds; specifically, we used 12 training and 4 testing bags per fold for the 16-bag track, and 16 training and 4 testing bags per fold for the 20-bag tracks. Thresholds were set as two standard deviations above the means of our heuristics $|v_{x}-\hat{v}_{x}|$ and $|\omega_{\psi}-\hat{\omega}_{\psi}|$. This process ensured that every bag was used exactly once for testing and 4 times for training within its respective track. The testing data from all folds were then used to evaluate slip detection accuracy, detection delay, and friction estimation accuracy across all tracks and surface conditions.
V-C Ground-Truth Friction Coefficient Measurement